Workshop on nonlinear dispersive and wave equations

Organized by Kagurazaka Analysis Seminer

日時、場所：１１月２４日（金曜日）１３：３０−１７：００理窓会館３階第２会議室日時、場所：１１月２５日（土曜日）１０：３０−１７：００東京理科大学１２号館１階 Y ゼミ１教室

１１月２４日と１１月２５日は会場がことなりますのでご注意ください。

平成１２年１１月２４日（金曜日）

13:30-14:10 Pavel I. Naumkin (Universidad Michoacana)

On the quadratic nonlinear Schrödinger equation

14:20-15:00 Tohru Ozawa (Hokkaido University)

Small data scattering for nonlinear Schrödinger equations

15:30-16:10 Masahito Ohta (Shizuoka University)

Instability of stationary states for nonlinear Schrödinger equations with magnetic field

16:20-17:00 Noboru Okazawa (Science University of Tokyo)

Evolution equations with simplest

nonlinear Ornstein-Uhlenbeck operators

平成１２年１１月２５日（土曜日）

10:30-11:10 Elena I. Kaikina (Instituto Technologico de Morelia) Large time behavior of solutions of the Korteweg-de Vries-Burgers equation on half-line 11:20-12:00 Hideo Kubo (Shizuoka University)

Global existence to wave equations with quadratic trems in two space dimensions

13:30-14:10 Kazuyoshi Yokoyama (Hokkaido Institute of Technology) Global existence of classical solutions for nonlinear wave equations in three space dimensions

14:20-15:00 Takeo Ushijima (Hokkaido University)

Blow-up problem to nonlinear heat equation with a nonlocal convection term

15:30-16:10 Kenji Nakanishi (Kobe University)

Nonrelativistic limit of Cauchy and scattering problems for nonlinear Klein-Gordon equations

16:20-17:00 Ryo Ikehata (Hiroshima University )

Decay of solutions for the dissipative wave equations in an exterior domain

本研究会は以下の助成によって行われます。文部省科学研究費基盤 B （２）課題番号 １２４４００５０「非線形分散型方程式の解の性質についての研究」